# Standard Lyndon loop words: weighted orders

### v1

@@ 1.1. Summary
for any total order on $\Delta^{+}$. Furthermore, a triangular decomposition
@@
(1.8) $\not\exists\ \alpha^{\prime},\beta^{\prime}\in\Delta^{+}\quad\text{s.t.}\quad\alpha<\alpha^{\prime}<\beta^{\prime}<\beta\quad\text{and}\quad\alpha+\beta=\alpha^{\prime}+\beta^{\prime}\,,$
@@
established in [11]. This induces the lexicographical order on $\Delta^{+}$ via
@@
is convex, cf. Proposition 3.18. As such, this order comes from a certain reduced word in the affine Weyl group associated to ${\mathfrak{g}}$ (= the Coxeter group associated to $\widehat{{\mathfrak{g}}}$), in accordance with Theorem 4.7. Therefore, the root vectors (1.11) exactly match (up to constants) the classical construction of [2, 4, 15], once we pass it through the “affine to loop” isomorphism of Theorem 5.14.
@@ 1.3. Acknowledgement
A.T. is deeply indebted to Andrei Neguţ for sharing his invaluable insights, teaching the Lyndon word’s theory, and stimulating discussions through the entire project. A.T. is grateful to INdAM-GNSAGA and FAR UNIMORE project CUP-E93C2300204000 for the support and wonderful working conditions during his visit to Italy in June 2024, where the final version of this note was prepared.
@@ 2.1. Lyndon words
Let $I$ be a finite ordered alphabet, and let $I^{*}$ be the set of all finite length words in the alphabet $I$. For $u=[i_{1}\dots i_{k}]\in I^{*}$, we define its length by $|u|=k$. We introduce the lexicographical order on $I^{*}$ in a standard way:
@@
The factorization (2.1) is called a costandard factorization of a Lyndon word.
@@
The factorization (2.2) is called a canonical factorization of a word.
@@ 2.7. Standard Lyndon words
    For a word $w\in I^{*}$ with a canonical factorization $w=\ell_{1}\dots\ell_{k}$ of (2.2), we set
@@
(2.6) $U({\mathfrak{a}})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon words}}{\mathbb{Q}}\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\,=\\ \bigoplus_{w\text{--standard words}}{\mathbb{Q}}\cdot e_{w}\ =\bigoplus_{w\text{--standard words}}{\mathbb{Q}}\cdot\,_{w}e\,.$
@@ 2.14. Application to simple Lie algebras
(2.11) $\deg\,[i_{1}\dots i_{k}]=\alpha_{i_{1}}+\dots+\alpha_{i_{k}}\in Q^{+}\,.$
@@ 2.15. Results of Leclerc
(2.13) $\alpha<\alpha+\beta<\beta$
@@
Consider the order on $\Delta^{+}$ induced from the lexicographical order on standard Lyndon words:
(2.14) $\alpha<\beta\quad\Longleftrightarrow\quad\ell(\alpha)<\ell(\beta)\ \ \mathrm{lexicographically}\,.$
@@
3. Loop standard Lyndon words
@@
We shall now introduce a family of total orders on ${\mathcal{I}}$, which will thus induce lexicographical orderings on loop words (3.1). To this end, we fix a total order on $I$ and choose a tuple of positive integers $\{c_{i}\}_{i\in I}\in{\mathbb{Z}}_{>0}^{I}$ (we call $c_{i}$ the weight of $i$). Following [16, Remark 3.15], we shall compare the loop letters of ${\mathcal{I}}$ via (1.10):
@@
(3.3) ${\mathcal{I}}^{(s)}=\left\{i^{(d)}\,\Big{|}\,i\in I,-s\cdot c_{i}\leq d\leq s\cdot c_{i}\right\}\qquad\forall\,s\in{\mathbb{Z}}_{\geq 0}\,.$
@@ 3.1. Filtration and basic properties
We now wish to extend Definition 2.10 in order to obtain a notion of standard (Lyndon) loop words, but here we must be careful as the alphabet ${\mathcal{I}}$ is infinite. In particular, the key assumption “for any word $v$, there are only finitely many words $u$ of the same length and $>v$ in the lexicographical order” of [11, §2] clearly fails. To deal with this issue, we consider the increasing filtration:
@@
(3.4) $L{\mathfrak{n}}^{+}\supset L^{(s)}{\mathfrak{n}}^{+}=\bigoplus_{\alpha\in\Delta^{+}}\bigoplus_{d=-s\cdot f(\alpha)}^{s\cdot f(\alpha)}{\mathbb{Q}}\cdot e_{\alpha}^{(d)}\qquad\forall\,s\in{\mathbb{Z}}_{\geq 0}\,.$
@@
The proof if completely analogous to that of [16, Proposition 2.25]. ∎
@@ 3.4. Exponent tightness
Let $v=[i_{1}^{(k_{1})}\dots\,i_{m}^{(k_{m})}]$ and $w=[j_{1}^{(t_{1})}\dots\,j_{m}^{(t_{m})}]$ be two exponent-tight loop words such that hdeg $\text{hdeg }w=\text{hdeg }v$ and vdeg $\text{vdeg }w=\text{vdeg }v+1$. Then:
(a) The first letter $j_{1}^{(t_{1})}$ of the loop word $w$ equals $\max_{1\leq a\leq m}\,\{i_{a}^{(d_{a}+1)}\}$.
(b) The multisets of the other letters coincide: $\{i_{a}^{(k_{a})}\}_{a=1}^{m}-\{j_{1}^{(t_{1}-1)}\}=\{j_{a}^{(t_{a})}\}_{a=2}^{m}$.
@@
Let $i_{r}^{(d_{r}+1)}=\max_{1\leq a\leq m}\,\{i_{a}^{(d_{a}+1)}\}$. Since $v$ is exponent-tight, so is any loop word $u$ formed by the letters $\{i_{a}^{(d_{a})}\}_{a\neq r}\cup\{i_{r}^{(d_{r}+1)}\}$ (a loop word is exponent-tight iff any loop word formed by the same multiset of letters is exponent-tight). But then $w$ and $u$ must have the same multisets of letters, according to Lemma 3.7. Since the loop word $u$ satisfies both properties (a) and (b), so does $w$. ∎
@@
Following the setup of Proposition 3.8, one may vice-versa express the multiset of letters of $v$ through the one for $w$: $\{i_{a}^{(k_{a})}\}_{a=1}^{m}=\{j_{1}^{(t_{1}-1)}\}\cup\{j_{a}^{(t_{a})}\}_{a=2}^{m}$.
@@
The base case of the induction is $n=2$. Let $\ell(\alpha,d)=[i_{1}^{(k_{1})}i_{2}^{(k_{2})}]$, where $i_{1}^{(k_{1})}<i_{2}^{(k_{2})}$ and $i_{1}\neq i_{2}$. We claim that $i_{1}^{(k_{1}-1)}>i_{2}^{(k_{2}+1)}$, as otherwise we would get $\ell(\alpha,d)=[i_{1}^{(k_{1})}i_{2}^{(k_{2})}]<[i_{1}^{(k_{1}-1)}i_{2}^{(k_{2}+1)}]$, a contradiction with Leclerc’s algorithm (3.6). But then, invoking (3.6), we obtain $\ell(\alpha,d)\geq i_{2}^{(k_{2}+1)}i_{1}^{(k_{1}-1)}$. This implies the desired inequality $i_{1}^{(k_{1})}\geq i_{2}^{(k_{2}+1)}$, establishing the base of the induction.
@@
with $f(\alpha)$ defined in (3.2). In particular, if $c_{i}=N>1$ for all $i\in I$ (thus the order on ${\mathcal{I}}$ is the same as for $c_{i}=1$ and so $\ell(\alpha,d)$ are the same as in [16]), then (3.12) only implies $|k_{i}-k_{j}|\leq N$, while Theorem 3.6 implies a much finer bound $|k_{i}-k_{j}|\leq 1$.
@@
Let $\ell(\alpha,d)=[i_{1}^{(k_{1})}\dots\,i_{n}^{(k_{n})}]$. Then $i_{1}^{(k_{1})}\leq i_{r}^{(k_{r})}$ and so $\frac{k_{1}}{c_{1}}\geq\frac{k_{r}}{c_{r}}$ for any $r$. Thus if $k_{1}\leq 0$, then $k_{r}\leq 0$ for any $r$, and so $d=\sum_{r=1}^{n}k_{r}\leq 0$, implying part (a).
To prove (b), we note that $i_{1}^{(k_{1})}\geq i_{r}^{(k_{r}+1)}$ for any $r$ by Theorem 3.6, hence $\frac{k_{1}}{c_{1}}\leq\frac{k_{r}+1}{c_{r}}$. If $k_{1}>0$, then $k_{r}\geq 0$ for all $r$, and so $d=\sum_{r=1}^{n}k_{r}>0$, a contradiction. ∎
@@ 3.14. Periodicity
(3.16) $U(L{\mathfrak{n}}^{+})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon loop words}}{\mathbb{Q}}\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\,=\\ \mathop{\bigoplus}_{w\text{--standard loop words}}{\mathbb{Q}}\cdot e_{w}\ =\mathop{\bigoplus}_{w\text{--standard loop words}}{\mathbb{Q}}\cdot\,_{w}e\,.$
@@ 3.17. Convexity and minimality
(3.17) $\ell(\alpha,d)<\ell(\alpha+\beta,d+e)<\ell(\beta,e)$
@@ 4. Lyndon words and Weyl groups
In this Section, we show that the lexicographic order (1.13) on $\Delta^{+}\times{\mathbb{Z}}$ induced by (3.13) is related to the construction of [19, 20] applied to a reduced decomposition of a translation element in the extended affine Weyl group encoding the weights $c_{i}$.
@@ 4.1. Affine Lie algebras
In this Section, we recall the next simplest class of Kac-Moody Lie algebras after the simple ones, the affine Lie algebras. Let ${\mathfrak{g}}$ be a simple finite-dimensional Lie algebra, $\{\alpha_{i}\}_{i\in I}$ be the simple roots, and $\theta\in\Delta^{+}$ be the highest root. The labels of the Dynkin diagram of ${\mathfrak{g}}$ are the positive integers $\{\theta_{i}\}_{i\in I}$ such that
@@
As opposed from the non-degenerate pairing on ${\mathfrak{g}}$ itself, the pairing on affine type root systems has a one-dimensional kernel, which is spanned by the minimal imaginary root $\delta=\alpha_{0}+\theta=(0,1)\in Q\times{\mathbb{Z}}$. This implies the fact that:
@@
(4.4) $c=h_{0}+\sum_{i\in I}\theta_{i}h_{i}\qquad\mathrm{is\ a\ central\ element\ of}\ \ \widehat{{\mathfrak{g}}}\,.$
@@
(4.5) $\widehat{\Delta}^{+}=\big{\{}\Delta^{+}\times{\mathbb{Z}}_{\geq 0}\big{\}}\sqcup\big{\{}0\times{\mathbb{Z}}_{>0}\big{\}}\sqcup\big{\{}\Delta^{-}\times{\mathbb{Z}}_{>0}\big{\}},$
(4.6) $\widehat{\Delta}^{-}=\big{\{}\Delta^{-}\times{\mathbb{Z}}_{\leq 0}\big{\}}\sqcup\big{\{}0\times{\mathbb{Z}}_{<0}\big{\}}\sqcup\big{\{}\Delta^{+}\times{\mathbb{Z}}_{<0}\big{\}}.$
@@
(4.7) $\widehat{{\mathfrak{g}}}=\widehat{{\mathfrak{h}}}\oplus\bigoplus_{\alpha\in\widehat{\Delta}}\widehat{{\mathfrak{g}}}_{\alpha}\quad\mathrm{where}\quad\widehat{{\mathfrak{h}}}\subset\widehat{{\mathfrak{g}}}-\mathrm{Cartan\ subalgebra}\,.$

### v2

@@ 1.1. Summary
for any total order on $\Delta^{+}$, with ${\mathbb{N}}={\mathbb{Z}}_{\geq 0}$. Furthermore, a triangular decomposition
@@
[…]
@@
established in [11]. This induces the lexicographic order on $\Delta^{+}$ via
@@
is convex, cf. Proposition 3.18. In fact, this order comes from a certain reduced word in the affine Weyl group associated to ${\mathfrak{g}}$ (= the Coxeter group associated to $\widehat{{\mathfrak{g}}}$), in accordance with Theorem 4.7. Therefore, the root vectors (1.10) exactly match (up to constants) the classical construction of [2, 4, 15], once we pass it through the “affine to loop” isomorphism of Theorem 5.14.
@@ 1.2. Outline
    In Section 6, we adapt most of our results to more general orders (6.1) on ${\mathcal{I}}$.
  • $\bullet$
@@ 1.3. Acknowledgement
A.T. is deeply indebted to Andrei Neguţ for sharing his invaluable insights, teaching the Lyndon word’s theory, and stimulating discussions through the entire project. A.T. is grateful to INdAM-GNSAGA and FAR UNIMORE project CUP-E93C2300204000 for the support and wonderful working conditions during his visit to Italy in June 2024, where the final version of this note was prepared. We are very grateful to the referees for their useful suggestions that improved the exposition.
@@ 2.1. Lyndon words
Let $I$ be a finite ordered alphabet, and let $I^{*}$ be the set of all finite length words in the alphabet $I$. For $u=[i_{1}\dots i_{k}]\in I^{*}$, we define its length by $|u|=k$. We introduce the lexicographic order on $I^{*}$ in a standard way:
@@
The factorization (2.1) is called the costandard factorization of a Lyndon word.
@@
The factorization (2.2) is called the canonical factorization of a word.
@@ 2.7. Standard Lyndon words
    For a word $w\in I^{*}$ with the canonical factorization $w=\ell_{1}\dots\ell_{k}$ of (2.2), we set
@@
(2.6) $\begin{split}&U({\mathfrak{a}})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon words}}{\mathbb{Q}}\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\,=\\ &\qquad\qquad\qquad\bigoplus_{w\text{--standard words}}{\mathbb{Q}}\cdot e_{w}\ =\bigoplus_{w\text{--standard words}}{\mathbb{Q}}\cdot\,_{w}e\,.\end{split}$
@@ 2.14. Application to simple Lie algebras
$\deg\,[i_{1}\dots i_{k}]=\alpha_{i_{1}}+\dots+\alpha_{i_{k}}\in Q^{+}\,.$
@@ 2.15. Results of Leclerc
$\alpha<\alpha+\beta<\beta$
@@
Consider the order on $\Delta^{+}$ induced from the lexicographic order on standard Lyndon words:
$\alpha<\beta\quad\Longleftrightarrow\quad\ell(\alpha)<\ell(\beta)\ \ \mathrm{lexicographically}\,.$
@@
3. Standard Lyndon loop words
@@
We shall now introduce a family of total orders on ${\mathcal{I}}$, which will thus induce lexicographic orderings on loop words (3.1). To this end, we fix a total order on $I$ and choose a tuple of positive integers $\{c_{i}\}_{i\in I}\in{\mathbb{Z}}_{>0}^{I}$ (we call $c_{i}$ the weight of $i$). Following [16, Remark 3.15], we shall compare the loop letters of ${\mathcal{I}}$ via (1.9):
@@
(3.3) ${\mathcal{I}}^{(s)}=\left\{i^{(d)}\,\Big{|}\,i\in I,-s\cdot c_{i}\leq d\leq s\cdot c_{i}\right\}\qquad\forall\,s\in{\mathbb{N}}\,.$
@@ 3.1. Filtration and basic properties
We now wish to extend Definition 2.10 in order to obtain a notion of standard (Lyndon) loop words, but here we must be careful as the alphabet ${\mathcal{I}}$ is infinite. In particular, the key assumption “for any word $v$, there are only finitely many words $u$ of the same length and $>v$ in the lexicographic order” of [11, §2] clearly fails. To deal with this issue, we consider the increasing filtration:
@@
(3.4) $L{\mathfrak{n}}^{+}\supset L^{(s)}{\mathfrak{n}}^{+}=\bigoplus_{\alpha\in\Delta^{+}}\bigoplus_{d=-s\cdot f(\alpha)}^{s\cdot f(\alpha)}{\mathbb{Q}}\cdot e_{\alpha}^{(d)}\qquad\forall\,s\in{\mathbb{N}}\,.$
@@
The proof is completely analogous to that of [16, Proposition 2.25]. ∎
@@ 3.4. Exponent tightness
Let $v=[i_{1}^{(d_{1})}\dots\,i_{m}^{(d_{m})}]$ and $w=[j_{1}^{(t_{1})}\dots\,j_{m}^{(t_{m})}]$ be two exponent-tight loop words such that hdeg $\text{hdeg }w=\text{hdeg }v$, vdeg $\text{vdeg }w=\text{vdeg }v+1$, and $j_{1}^{(t_{1})}\leq j_{r}^{(t_{r})}$ for all $r$. Then:
(a) The first letter $j_{1}^{(t_{1})}$ of the loop word $w$ equals $\max_{1\leq a\leq m}\,\{i_{a}^{(d_{a}+1)}\}$;
(b) The multisets of the other letters coincide: $\{i_{a}^{(d_{a})}\}_{a=1}^{m}-\{j_{1}^{(t_{1}-1)}\}=\{j_{a}^{(t_{a})}\}_{a=2}^{m}$.
@@
Let $i_{r}^{(d_{r}+1)}=\max_{1\leq a\leq m}\,\{i_{a}^{(d_{a}+1)}\}$. Since $v$ is exponent-tight, so is any loop word $u$ formed by the letters $\{i_{a}^{(d_{a})}\}_{a\neq r}\cup\{i_{r}^{(d_{r}+1)}\}$ (a loop word is exponent-tight iff any loop word formed by the same multiset of letters is exponent-tight). But then $w$ and $u$ must have the same multisets of letters, according to Lemma 3.7. Since the loop word $u$ satisfies (b), $v$ is exponent-tight and $w$ starts with its smallest letter, we obtain both properties (a) and (b). ∎
@@
Following the setup of Proposition 3.8, one may vice-versa express the multiset of letters of $v$ through the one for $w$: $\{i_{a}^{(d_{a})}\}_{a=1}^{m}=\{j_{1}^{(t_{1}-1)}\}\cup\{j_{a}^{(t_{a})}\}_{a=2}^{m}$.
@@
The base case of the induction is $n=2$. Let $\ell(\alpha,d)=[i_{1}^{(d_{1})}i_{2}^{(d_{2})}]$, where $i_{1}^{(d_{1})}<i_{2}^{(d_{2})}$ and $i_{1}\neq i_{2}$. We claim that $i_{1}^{(d_{1}-1)}>i_{2}^{(d_{2}+1)}$, as otherwise we would get $\ell(\alpha,d)=[i_{1}^{(d_{1})}i_{2}^{(d_{2})}]<[i_{1}^{(d_{1}-1)}i_{2}^{(d_{2}+1)}]$, a contradiction with Leclerc’s algorithm (3.6). But then, invoking (3.6), we obtain $\ell(\alpha,d)\geq i_{2}^{(d_{2}+1)}i_{1}^{(d_{1}-1)}$. This implies the desired inequality $i_{1}^{(d_{1})}\geq i_{2}^{(d_{2}+1)}$, establishing the base of the induction.
@@
with $f(\alpha)$ defined in (3.2). In particular, if $c_{i}=N>1$ for all $i\in I$ (thus the order on ${\mathcal{I}}$ is the same as for $c_{i}=1$ and so $\ell(\alpha,d)$ are the same as in [16]), then (3.12) only implies $|d_{r}-d_{t}|\leq N$, while Theorem 3.6 implies a much finer bound $|d_{r}-d_{t}|\leq 1$.
@@
Let $\ell(\alpha,d)=[i_{1}^{(d_{1})}\dots\,i_{n}^{(d_{n})}]$. Then $i_{1}^{(d_{1})}\leq i_{r}^{(d_{r})}$ and so $\frac{d_{1}}{c_{i_{1}}}\geq\frac{d_{r}}{c_{i_{r}}}$ for any $r$. Thus if $d_{1}\leq 0$, then $d_{r}\leq 0$ for any $r$, and so $d=\sum_{r=1}^{n}d_{r}\leq 0$, implying part (a).
To prove (b), we note that $i_{1}^{(d_{1})}\geq i_{r}^{(d_{r}+1)}$ for any $r$ by Theorem 3.6, thus $\frac{d_{1}}{c_{i_{1}}}\leq\frac{d_{r}+1}{c_{i_{r}}}$. If $d_{1}>0$, then $d_{r}\geq 0$ for all $r$, and so $d=\sum_{r=1}^{n}d_{r}>0$, a contradiction. ∎
@@ 3.14. Periodicity
(3.16) $\begin{split}&U(L{\mathfrak{n}}^{+})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon loop words}}{\mathbb{Q}}\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\,=\\ &\qquad\qquad\qquad\mathop{\bigoplus}_{w\text{--standard loop words}}{\mathbb{Q}}\cdot e_{w}\ =\mathop{\bigoplus}_{w\text{--standard loop words}}{\mathbb{Q}}\cdot\,_{w}e\,.\end{split}$
@@ 3.17. Convexity and minimality
$\ell(\alpha,d)<\ell(\alpha+\beta,d+e)<\ell(\beta,e)$
@@ 4. Lyndon words and Weyl groups
In this Section, we show that the order (1.12) on $\Delta^{+}\times{\mathbb{Z}}$ induced by (3.13) is related to the construction of [19, 20] applied to a reduced decomposition of a translation element in the extended affine Weyl group encoding the weights $c_{i}$.
@@ 4.1. Affine Lie algebras
In this Subsection, we recall the next simplest class of Kac-Moody Lie algebras after the simple ones, the affine Lie algebras. Let ${\mathfrak{g}}$ be a simple finite-dimensional Lie algebra, $\{\alpha_{i}\}_{i\in I}$ be the simple roots, and $\theta\in\Delta^{+}$ be the highest root. The labels of the Dynkin diagram of ${\mathfrak{g}}$ are the positive integers $\{\theta_{i}\}_{i\in I}$ such that
@@
As opposed from the non-degenerate pairing on $Q$ itself, the pairing on affine type root systems has a one-dimensional kernel, which is spanned by the minimal imaginary root $\delta=\alpha_{0}+\theta=(0,1)\in Q\times{\mathbb{Z}}$. This implies the fact that:
@@
$c=h_{0}+\sum_{i\in I}\theta_{i}h_{i}\qquad\mathrm{is\ a\ central\ element\ of}\ \ \widehat{{\mathfrak{g}}}\,.$
@@
$\widehat{\Delta}^{+}=\big{\{}\Delta^{+}\times{\mathbb{Z}}_{\geq 0}\big{\}}\sqcup\big{\{}0\times{\mathbb{Z}}_{>0}\big{\}}\sqcup\big{\{}\Delta^{-}\times{\mathbb{Z}}_{>0}\big{\}},$
$\widehat{\Delta}^{-}=\big{\{}\Delta^{-}\times{\mathbb{Z}}_{\leq 0}\big{\}}\sqcup\big{\{}0\times{\mathbb{Z}}_{<0}\big{\}}\sqcup\big{\{}\Delta^{+}\times{\mathbb{Z}}_{<0}\big{\}}.$
@@
$\widehat{{\mathfrak{g}}}=\widehat{{\mathfrak{h}}}\oplus\bigoplus_{\alpha\in\widehat{\Delta}}\widehat{{\mathfrak{g}}}_{\alpha}\quad\mathrm{where}\quad\widehat{{\mathfrak{h}}}\subset\widehat{{\mathfrak{g}}}-\mathrm{Cartan\ subalgebra}\,.$